OpenStudy (help_people):

If h(x) = 5x − 3 and j(x) = −2x, solve h[j(2)] and select the correct answer below. 17 23 −17 −23

OpenStudy (danjs):

For a compound function like that,, figure the innermost one out first, then use that value and move outwards

OpenStudy (danjs):

what is j(2) = ?

OpenStudy (help_people):

to be honest i have no clue @DanJS

OpenStudy (help_people):

-4? or 4?

OpenStudy (danjs):

j(x) = -2x j(2) = -2 * 2 = -4

OpenStudy (danjs):

Now to calculate h[j(2)] ... you use that value for j(2), for x in h

OpenStudy (danjs):

h( - 4) = ...

OpenStudy (help_people):

Yay! I got it right (Sort of) and how would you find h(-4)

OpenStudy (help_people):

^ @DanJS

OpenStudy (danjs):

if h(x) = 5x-3 then h(-4) = 5*(-4) -3 = -20 - 3 = -23

OpenStudy (danjs):

for those, work the inner function first, then keep using those output values and work outwards to the next function like we did

OpenStudy (help_people):

ok so am i done or i shtere more to go ?

OpenStudy (help_people):

@DanJS

OpenStudy (danjs):

that is it, you figured the inner function out, which was -4, then you use that value in the outer function h(-4) that is it

OpenStudy (help_people):

sp the answer is -23 ? @DanJS

OpenStudy (help_people):

*so

OpenStudy (help_people):

and if it is ty so much :) @DanJS

OpenStudy (danjs):

yes